# Posterior Consistency in Parametric Models via a Tighter Notion of Identifiability

### v1

@@ Abstract
We investigate Bayesian posterior consistency in the context of parametric models with proper priors. While typical state-of-the-art approaches rely on regularity conditions that are difficult to verify and often misaligned with the actual mechanisms driving posterior consistency, we propose an alternative framework centered on a simple yet general condition we call sequential identifiability. This concept strengthens the usual identifiability assumption by requiring that any sequence of parameters whose induced distributions converge to the true data-generating distribution must itself converge to the true parameter value. We demonstrate that sequential identifiability, combined with a standard Kullback–Leibler prior support condition, is sufficient to ensure posterior consistency. Moreover, we show that failure of this condition necessarily entails a specific and pathological form of oscillations of the model around the true density, which cannot exist without intentional design. This leads to the important insight that posterior inconsistency may be safely ruled out, except in the unrealistic scenario where the modeler possesses precise knowledge of the data-generating distribution and deliberately incorporates oscillatory pathologies into the model targeting the corresponding density. Taken together, these results provide a unified perspective on both consistency and inconsistency in parametric settings, significantly expanding the class of models for which posterior consistency can be rigorously established. To illustrate the strength and versatility of our framework, we construct a one-dimensional model that violates standard regularity conditions and fails to admit a consistent maximum likelihood estimator, yet supports a complete posterior consistency analysis via our approach.
@@ 1 Introduction
Asymptotic consistency is a fundamental criterion for evaluating the quality of statistical estimation procedures. In Bayesian inference, the study of consistency of posterior distributions has been a highly active area of research, with substantial recent progress particularly in nonparametric settings. A prominent example is the ubiquitous problem of density estimation from independently and identically distributed (iid) observations, where key contributions by Ghosal et al., (1999), Barron et al., (1999) and Walker, (2004) address consistency by sieve methods ensuring that the posterior distribution concentrates in any appropriately defined neighborhood of the true data-generating density.
@@
In contrast, this article focuses on the study of posterior consistency in parametric models. To set the stage, consider the following standard modeling framework: let the sample space be $(\mathbb{R},\mathscr{B}(\mathbb{R}))$, where $\mathscr{B}(T)$ denotes the Borel $\sigma$-algebra on a topological space $T$, and let $\mathrm{d}x$ denote the Lebesgue measure on the real line. Define the statistical model $\mathcal{F}_{\Theta}:=\{f_{\theta}:\theta\in\Theta\}$, where $f_{\theta}:=\mathrm{d}F_{\theta}/\mathrm{d}x$ is the density (Radon–Nikodym derivative) associated with a probability measure [^1] With a slight abuse of notation, we identify any probability measure $F$ on $(\mathbb{R},\mathscr{B}(\mathbb{R}))$ with its cumulative disribution function (cdf), writing $F(x)\equiv F((-\infty,x])$ for all $x\in\mathbb{R}$. $F_{\theta}(\mathrm{d}x)\ll\mathrm{d}x$ parametrized by $\theta\in\Theta\subseteq\mathbb{R}^{p}$, for some $p\in\mathbb{N}$, and where $\Theta$ is assumed to be closed. As is natural for any sensibly parametrized family $\mathcal{F}_{\Theta}$ intended for density estimation, we assume that convergence of parameters in $\Theta$ is equivalent to convergence of the associated densities with respect to some strong metric, such as the Hellinger distance $d_{h}$; that is, for all $\theta,\theta_{1},\theta_{2},\ldots\in\Theta$,
@@
see Section 2 for a formal definition of $d_{h}$.
@@
Posterior consistency in parametric models was first studied in the foundational work of doob1949application, who showed that, under a mild asymptotic estimability condition, the posterior distribution is consistent at any $\theta_{\star}$ belonging to a set $\Theta_{\star}\subseteq\Theta$ of prior mass one. While remarkably parsimonious in its assumptions, Doob’s approach suffers from a major drawback: the set $\Theta_{\star}$ is not explicitly characterized by the theorem, making it impossible to verify whether a given $\theta_{\star}\in\Theta$ belongs to it. Moreover, the complement of $\Theta_{\star}$—where posterior inconsistency may occur—could, in principle, be large.
In contrast to Doob’s approach, the conditions currently considered state of the art for establishing parametric consistency rely on directly imposing smoothness and regularity assumptions on the model $\mathcal{F}_{\Theta}$. The work of Walker, (1969) (see also Berk,, 1970) is a prominent example that continues to be cited in contemporary applied and methodological literature on Bayesian methods (e.g., Mao et al.,, 2024; Rustand et al.,, 2023; Miller,, 2021; Doğan et al.,, 2021). Specifically, Walker, (1969) established posterior consistency by requiring the log-likelihood mapping $\theta\mapsto\ln f_{\theta}(x)$ to possess a sufficiently regular modulus of continuity and to exhibit suitable decay outside compact subsets of $\Theta$ (see Section 3 for a detailed discussion). Under these conditions, consistency of both the maximum likelihood estimator (MLE) and the posterior distribution follows. Although conditions of this type remain the gold standard for establishing parametric consistency (see, e.g., the recent textbook treatment in Ghosal and van der Vaart,, 2017, sect. 6.4), in this article we challenge their suitability by demonstrating that they are overly restrictive, misaligned with the fundamental mechanisms governing posterior convergence, and often difficult to verify in parametric models beyond the most basic cases.
1.1 Our contribution
To address this gap and develop a unified framework for the study of parametric posterior consistency, we advocate a foundational reexamination of the problem. Our starting point is the seminal result of Schwartz, (1964), who showed that if the model is identifiable and $\theta_{\star}$ lies in the Kullback–Leibler (KL) support of the prior $\Pi$ (see Section 2 for a precise definition), then the posterior is weakly consistent at $\theta_{\star}$, i.e., consistent with respect to the topology of weak convergence. This is equivalent to stating that the posterior concentrates on parameter values $\theta\in\Theta$ whose associated cdfs $F_{\theta}$ are close to $F_{\theta_{\star}}$, or, more precisely, those $F_{\theta}$ lying in any arbitrarily small weak neighborhood of $F_{\theta_{\star}}$.
The difficulty in translating weak consistency into parameter consistency lies in the fact that the standard assumption of identifiability does not account for weak limiting behavior. As a result, it remains possible for a sequence of cdfs in the model to converge to $F_{\theta_{\star}}$ while the corresponding sequence of parameters fails to converge to $\theta_{\star}$. This creates a gap between weak consistency and consistency in the sense of parameter values: under the assumptions of Schwartz’s theorem, the posterior may accumulate on regions of the parameter space that yield the correct distribution function but are not themselves close to the true parameter value. To address this loophole, our key contribution is to introduce a new notion of identifiability, which we term sequential identifiability, requiring that no sequence of parameters failing to converge to $\theta_{\star}$ can give rise to a sequence of cdfs converging to $F_{\theta_{\star}}$. In light of Schwartz’s theorem, and by incorporating weak limits into its formulation, sequential identifiability ensures that the posterior concentrates around the true parameter value and, consequently, the true density function.
We argue that, in parametric settings, sequential identifiability provides a fundamental yet non-invasive tightening of standard identifiability. Indeed, we show that it holds for many widely used parametric models and that violations leading to inconsistency can occur only when the model includes densities that oscillate arbitrarily often around the true density. Consequently, unless the modeler possesses precise knowledge of the data-generating process and deliberately incorporates specifically designed pathological oscillations—two highly unrealistic conditions in any practical parametric modeling scenario—the model can be declared sequentially identifiable and, by default, consistent. To illustrate these ideas, we construct a parametric model based on cosine oscillations that violates classical regularity conditions and fails to satisfy MLE consistency, yet exhibits posterior consistency at all but one parameter value, thanks to sequential identifiability. The extreme oscillatory behavior of this model, together with the fact that issues with sequential identifiability arise only at a single parameter value, illustrates our point: failures of posterior consistency can be almost universally ruled out in parametric models of practical relevance, unless one intentionally designs a model that violates sequential identifiability by knowing which true density to target with oscillations.
1.2 Summary of the article
@@
The rest of the article is structured as follows. After establishing basic notation and definitions in Section 2, Section 3 reviews some classical conditions proposed in the literature to address posterior consistency in parametric models. In Section 4, we introduce our alternative approach based on sequential identifiability, providing general sufficient conditions for posterior consistency and linking inconsistency to a peculiar oscillatory behavior of the densities in the model. Section 5 illustrates our framework through a one-dimensional parametric model which, despite violating classical regularity conditions, is readily handled using our methodology. This example also serves to highlight key aspects of our theoretical analysis, particularly in relation to oscillations. Section 6 concludes the paper, while the proofs of all the theoretical results can be found in the supplementary material at the end of the article.
@@ 3 Classical conditions for parametric posterior consistency
Intuitively, assumption W1 ensures that the support of the model remains fixed across parameter values, thereby preventing singularities in likelihood ratios. Assumption W2 imposes a form of local uniform continuity on the log-likelihood function with respect to the parameter, captured via a modulus of continuity that vanishes both pointwise and in expectation under any true model. Finally, assumption W3 provides control over the relative tail behavior of the log-likelihood, requiring that log-likelihood ratios decay sufficiently fast—in an integrable manner—as the parameter value diverges.
Walker, (1969) showed that assumptions W1–W3, together with a positive and continuous prior density, imply posterior consistency at any $\theta_{\star}\in\Theta$ by ensuring MLE consistency and controlling the likelihood ratio away from the true parameter value. While technically valid—and while the positivity of the prior density may, under mild regularity conditions, be viewed as equivalent to the KL support condition of Schwartz, (1964)—these assumptions are quite specific and restrictive (see, e.g., Section 5 for a demonstration of this point). More importantly, their formulation is not informed by any clear connection to the core mechanisms that govern posterior consistency. Before presenting our own alternative treatment, we briefly review another classical approach to parametric consistency which, although potentially more broadly applicable, shares many of the same underlying limitations.

### v2

@@ Abstract
We investigate Bayesian posterior consistency in parametric density models with proper priors, challenging the common perception that the topic is settled. Classical results from the 1970s established posterior consistency as a consequence of MLE convergence, by combining regularity conditions with the assumption of model identifiability. In particular, the latter was treated as a background assumption and never examined in depth. This approach has gone largely unquestioned, partly due to a subsequent and nearly exclusive shift in focus to sieve-based methods tailored to nonparametric consistency. In our analysis, we place identifiability at the heart of posterior consistency. We show that, once one enlarges the model family to include weak limits, inconsistency fundamentally stems from a failure of identifiability at the true distribution. This finding reveals an important distinction: while such a failure can occur naturally in nonparametric models, it is highly implausible and essentially self-inflicted in parametric ones. This motivates a separate treatment of the two cases, with our focus here on the parametric side. Our theory leads to the finding that classical regularity assumptions are overly restrictive, while a simple tightening of identifiability suffices to establish posterior consistency even in irregular models where the MLE is inconsistent. Moreover, we prove that inconsistency requires the presence of densities with pathological oscillations that precisely match the true distribution. As we exemplify with an illustrative model, such behavior may arise only if the modeler possesses exact prior knowledge of the ground truth and adversarially encodes it in the model. Our example also underscores the need for distinct tools to study frequentist and Bayesian consistency: while MLE inconsistency stems from overfitting caused by likelihood peaks at the data—appropriately addressed through regularity or sieve methods—Bayesian parametric inconsistency is more naturally resolved by examining the identifiability structure of the model.
@@ 1 Introduction
Asymptotic consistency is a fundamental criterion for evaluating the quality of statistical estimation procedures. In Bayesian inference, the study of consistency of posterior distributions has been a highly active area of research, with the first contributions dating back to Joseph Doob’s work (Doob,, 1949). In his seminal result, Doob proved that, under a mild estimability condition, the posterior distribution asymptotically concentrates in arbitrary neighborhoods of the data-generating parameter value (under the assumption that a true one exists), for any parameter value belonging to a set of prior mass one. While remarkably parsimonious in its assumptions, Doob’s approach was later reevaluated due to a major drawback: the theorem does not explicitly characterize the set of parameters at which the posterior is consistent, making it impossible to determine whether a given parameter belongs to this set, or how large its complement (where inconsistency may occur) is.
Subsequent major contributions appeared in the late 1960s and through the 1970s, focusing on parametric models for density estimation based on independently and identically distributed (iid) observations. A prominent example is the work of Andrew Walker (Walker,, 1969) (see also Berk,, 1970), which leveraged the extensive classical theory on the consistency of frequentist procedures, particularly maximum likelihood estimators (MLEs), to establish posterior consistency. These approaches involved imposing identifiability and regularity conditions on the parametric family of likelihoods—such as smoothness of the log-likelihood ratio near zero and appropriate decay outside compact sets—to ensure both MLE consistency and concentration of the posterior around the MLE, thereby guaranteeing the convergence of posterior mass towards the true parameter value.
With Thomas Ferguson’s breakthrough definition of the Dirichlet process prior (Ferguson,, 1973), which made nonparametric inference practically feasible within the Bayesian framework, the literature on Bayesian consistency quickly shifted focus to infinite-dimensional models—partly due to the intriguing mathematical challenges they posed. Key early contributions include Diaconis and Freedman, 1986b ; Diaconis and Freedman, 1986a , followed by major developments in nonparametric density estimation (Ghosal et al.,, 1999; Barron et al.,, 1999; Walker,, 2004). The core idea in these works is to establish consistency using sieve methods which ensure that the posterior distribution concentrates within suitably defined neighborhoods of the true data-generating density. These approaches also paved the way for important theoretical advances—such as the study of contraction rates (Ghosal et al.,, 2000; Ghosal and van der Vaart, 2007a, ; Lijoi et al.,, 2007) and misspecified models (Kleijn and van der Vaart,, 2006; De Blasi and Walker,, 2013)—as well as for applications to specific nonparametric models (Lijoi et al.,, 2005; Ghosal,, 2001; Ghosal and van der Vaart, 2007b, ; Ghosal and Roy,, 2006).
This trajectory of the literature has effectively shifted focus away from the parametric setting, with the study of nonparametric consistency nearly halting progress on the parametric side. As a result, conditions dating back more than 50 years, such as those in Walker, (1969), remain the standard tools for establishing consistency in finite-dimensional models. [^1] See, e.g., the recent textbook treatment in Ghosal and van der Vaart, (2017), sect. 6.4. See instead Mao et al., (2024); Rustand et al., (2023); Miller, (2021); Doğan et al., (2021) for very recent applied and methodological work citing Walker, (1969). Indeed, to the best of our knowledge, although posterior consistency has been extensively studied for certain specific parametric models in recent years, [^2] For instance, in the case of finite mixtures (Rousseau and Mengersen,, 2011; Guha et al.,, 2021) there has been little to no major development in the realm of general parametric consistency since the original contributions from the early 1970s. In this article, we argue that this has resulted in a narrow view of parametric consistency and left key aspects of the topic underexplored. Moreover, these trends have limited the available conditions for establishing consistency to those developed decades ago—a time when, due to the lack of tools for a genuinely Bayesian analysis, the best available approach was to tie posterior consistency to the regular behavior of the MLE, thereby evaluating the asymptotic performance of Bayesian procedures through their alignment with frequentist ones.
1.1 A new understanding of parametric consistency
@@
The key aim of this article is to conduct a fundamental reexamination of posterior parametric consistency. To set the stage for our discussion, consider the following standard modeling framework: let the sample space be $(\mathbb{R},\mathscr{B}(\mathbb{R}))$, [^3] Throughout the article, we only consider $\mathbb{R}$ as our sample space to facilitate exposition, though much of our treatment easily extends to higher-dimensional scenarios. where $\mathscr{B}(T)$ denotes the Borel $\sigma$-algebra on a topological space $T$, and let $\mathrm{d}x$ denote the Lebesgue measure on the real line. Define the statistical model $\mathcal{F}_{\Theta}:=\{f_{\theta}:\theta\in\Theta\}$, where $f_{\theta}:=\mathrm{d}F_{\theta}/\mathrm{d}x$ is the density (Radon–Nikodym derivative) associated with a probability measure [^4] With a slight abuse of notation, we identify any probability measure $F$ on $(\mathbb{R},\mathscr{B}(\mathbb{R}))$ with its cumulative disribution function (CDF), writing $F(x)\equiv F((-\infty,x])$ for all $x\in\mathbb{R}$. $F_{\theta}(\mathrm{d}x)\ll\mathrm{d}x$ parametrized by $\theta\in\Theta\subseteq\mathbb{R}^{p}$, for some $p\in\mathbb{N}$, and where $\Theta$ is assumed to be closed. As is natural for any sensibly parametrized family $\mathcal{F}_{\Theta}$ intended for density estimation, we assume that convergence of parameters in $\Theta$ is equivalent to convergence of the associated densities with respect to some strong metric, such as the Hellinger distance $d_{h}$; that is, for all $\theta,\theta_{1},\theta_{2},\ldots\in\Theta$,
@@
see Section 2 for a formal definition of $d_{h}$. This allows us to identify $(\Theta,\|\cdot\|)$ with $(\mathcal{F}_{\Theta},d_{h})$ in terms of their metric structure, and to think of sequences of parameters $\theta_{1},\theta_{2},\ldots\in\Theta$ as sequences of densities in the model class $\mathcal{F}_{\Theta}$.
@@
With this modeling framework in mind, our starting point is the seminal result by Lorraine Schwartz (Schwartz,, 1964, see also Theorem 1 below), who proved that if $\theta_{\star}$ lies in the Kullback-Leibler (KL) support of $\Pi$ (see Definition 1 below), then the posterior distribution is weakly consistent at $F_{\theta_{\star}}$, that is, (1) holds when $A_{\varepsilon}$ is replaced with a neighborhood of $F_{\theta_{\star}}$ in the weak topology on $\mathcal{F}_{\Theta}$, metrized, for instance, by the Lévy-Prokhorov metric (see Section 2 for a formal definition). In other words, Schwartz’s theorem guarantees that, under a natural prior support condition, the posterior distribution concentrates in any small neighborhood of the true CDF. This result is remarkable in that it requires only a well-specified prior and applies even in nonparametric models—i.e., when $\Theta$ is infinite-dimensional rather than Euclidean. However, this type of consistency is not sufficient to characterize the posterior’s behavior with respect to neighborhoods of the true density, rather than of the true CDF. The issue is that the posterior may concentrate on regions of the parameter space that yield arbitrarily good approximations of the true distribution in the weak sense, while still being far from the true density itself. A more detailed exploration of this phenomenon is provided in Section 4, but intuition can already be gained from Figure 1, which shows how a sequence of densities with increasing oscillations can converge in distribution to a target CDF while failing to converge in any meaningful to a proper density function.
Figure 1: The initial segment of a sequence $f_{1},f_{2},\dots$ of indefinitely oscillating densities (solid gray, top row), whose corresponding sequence of CDFs (solid gray, bottom row) converges to a proper limiting distribution (dashed black). Such convergence is possible because the density sequence exhibits increasingly frequent oscillations around the density associated with the limiting CDF.
Interestingly, the strength of Schwartz’s result and of the proof techniques it introduced, based on exponential hypothesis tests, had a significant influence on subsequent works in nonparametric consistency. In particular, Schwartz’s approach was later recognized (Ghosal and van der Vaart,, 2017) as central to the sieve-based strategies of Barron et al., (1999); Ghosal et al., (1999), which rely on controlling the complexity (e.g., the $L^{1}$ metric entropy) of certain sets of densities on which most of the prior mass is placed. In contrast, Schwartz’s result has received little attention in the parametric literature, likely due to the perception that its full potential had already been realized in the nonparametric setting. As a consequence, consistency results for parametric models have either relied on the classical conditions from the 1970s, such as those of Walker, (1969), or have been treated as simple corollaries of more abstract nonparametric analyses involving tools like sieves and entropy bounds.
However, we contend that these historical developments were shaped by a flawed assumption: that Bayesian parametric and nonparametric models for density estimation can be meaningfully analyzed through the same lens. Our disagreement with this assumption stems from the following observation. Schwartz’s result implies that the posterior asymptotically concentrates on parameter values that yield the true distribution function. Therefore, inconsistency in terms of densities (or parameter values) can occur only if there are regions of the parameter space, away from the true parameter $\theta_{\star}$, that produce CDFs arbitrarily close to $F_{\theta_{\star}}$ (and the posterior happens to concentrate around those instead of $\theta_{\star}$ itself). In other words, once the parameter space is enlarged to include all weak limits of sequences of distributions in the model, inconsistency can arise only if the true CDF corresponds to at least two distinct regions of the parameter space: a “correct” one (around $\theta_{\star}$) and at least one additional “incorrect” one (away from $\theta_{\star}$) that is obtained as the weak limit of some sequence of CDFs in the model. This situation, then, is best understood as a problem of identifiability at $\theta_{\star}$ in the enlarged parameter space, and if one is able to rule out the lack of identifiability described above, then consistency in terms of CDFs, guaranteed by Schwartz’s theorem, directly implies consistency in terms of parameters.
Once the key issue in achieving posterior consistency is recognized to be this extended notion of identifiability, which we later term sequential identifiability (see Definition 3 below), a fundamental discrepancy between parametric and nonparametric models becomes apparent. Nonparametric models, by their very nature, are inherently prone to a lack of sequential identifiability over large portions of the parameter space. Consider, for example, infinite Gaussian mixtures, which can weakly approximate any continuous distribution function while failing to converge in density due to excessively large values of the precision parameter. In such cases, the assumption of sequential identifiability no longer holds unless the prior support is appropriately constrained, and a nuanced analysis is required to account for the oscillatory behavior that gives rise to unidentifiability (see Sections 4 and 6 for further elaboration).
On the other hand, the finite-dimensional nature of parametric models calls for a fundamentally different approach to sequential identifiability. For any parametric family that is identifiable in the usual sense, the inherently limited approximation power of the finite-dimensional parameter space—relative to the unbounded flexibility we will show to be required to induce sequential unidentifiability—makes it effectively impossible to weakly approximate more than a handful of distribution away from the corresponding parameter values. As we demonstrate in Section 5, even a highly oscillatory family of densities, like the one we design for illustration, may fail sequential identifiability at most at a few isolated parameter values (just a single one, in our example). More generally, Theorem 3 shows that for sequential unidentifiability to occur at the true parameter, the model must include densities with arbitrarily frequent, finely tuned oscillations that integrate exactly to the true distribution function. Such behavior is highly implausible in any practical modeling context: because a parametric model allows, at most, for a small number of sequentially unidentifiable parameter values, posterior inconsistency would require exact prior knowledge of the true density in order to engineer the oscillations to align precisely with one of these few values, making inconsistency an unrealistic concern when dealing with parametric families genuinely designed for statistical modeling. [^5] We note that Walker et al., (2005) identified the phenomenon of data tracking, arising from oscillatory densities, as a potential source of inconsistency. Our treatment, however, comprehensively reframes the issue in terms of identifiability and precisely characterizes the oscillatory patterns required for inconsistency to arise, showing that such behavior is not merely one possible cause, but in fact the sole mechanism (implausible in parametric models) through which inconsistency can occur.
To summarize, the key distinction between parametric and nonparametric models lies in how they satisfy Schwartz’s theorem. In nonparametric models, the posterior can concentrate on densities that flexibly oscillate around (nearly) any candidate data-generating process, allowing them to weakly approximate the true distribution even when far from the true density—in our terminology, sequential unidentifiability holds over the whole (or most of) the model space. In contrast, parametric models generally lack this flexibility, so for Schwartz’s result to apply, the posterior must concentrate at the true parameter value.
Our analysis also highlights how, in a certain sense, the classical literature on parametric posterior consistency followed the MLE consistency literature in the “wrong” direction. Indeed, the theory of MLE consistency begins with the standard assumption of identifiability and adds suitable regularity conditions to ensure convergence. Parametric posterior consistency was then made to follow MLE consistency down the path of regularity, [^6] One may argue that this was also true in the nonparametric setting, where the sieve conditions used to prove posterior consistency closely mirror those used for sieve MLE convergence (e.g., see Wong and Shen,, 1995; Shen and Wong,, 1994; van de Geer,, 2000). at a time when the literature had not yet recognized that, in light of Schwartz’s theorem, identifiability—not regularity—is the truly fundamental requirement. A central contribution of our work is to bring identifiability back to the core of the analysis of posterior consistency, yielding substantially better conditions in the parametric setting. Moreover, as the example in Section 5 illustrates, our strengthened notion of identifiability is specifically tailored to Bayesian procedures and has no bearing on MLE consistency, which may still fail due to the likelihood’s tendency to form peaks at the data and despite sequential identifiability at the data-generating parameter.
In summary, treating the parametric case separately from the nonparametric one, while still adopting Schwartz’s weak consistency result as a common foundation, naturally brings sequential identifiability to the forefront as the most natural and minimal condition for consistency in parametric models. As we show throughout the paper, this revised perspective allows us to go beyond the classical regularity assumptions from earlier works, enabling consistency even in irregular models where the maximum likelihood estimator performs poorly, such as our cosine-based example. This possibility reflects the different mechanisms behind inconsistency of frequentist and Bayesian procedures. For maximum likelihood, inconsistency often results from the likelihood overfitting the data, away from the true parameter value, by placing peaks at the observed points. Regularity or sieve conditions are appropriately designed to prevent this behavior, as seen in the classical assumptions discussed in Section 3, particularly assumption W3, as well as in our example in Section 5, where these conditions fail because the model’s oscillatory behavior allows each data point to be placed at a likelihood peak as the parameter diverges. By contrast, Schwartz’s theorem implies that Bayesian posteriors are naturally able to recover the true distribution function, making sequential identifiability alone sufficient to ensure posterior consistency even in cases where the MLE is inconsistent.
Layout of the paper.
@@
The rest of the article is structured as follows. After establishing basic notation and definitions in Section 2, Section 3 reviews in detail the conditions proposed in the classical literature on posterior parametric consistency. In Section 4, we formally introduce our foundational approach based on sequential identifiability, providing general sufficient conditions for posterior consistency and linking inconsistency to a peculiar oscillatory behavior of the densities in the model. Section 5 illustrates our framework through a one-dimensional parametric model which, despite violating classical regularity assumptions, is readily handled using our methodology. This example also serves to highlight key aspects of our theoretical analysis, particularly in relation to oscillations. Section 6 concludes the paper, while the proofs of all the theoretical results can be found in the supplementary material at the end of the article.
@@ 3 Classical conditions for parametric posterior consistency
Intuitively, assumption W1 ensures that the support of the model remains fixed across parameter values, thereby preventing singularities in likelihood ratios. Assumption W2 imposes a form of local uniform continuity on the log-likelihood function with respect to the parameter, captured via a modulus of continuity that vanishes both pointwise and in expectation under any true model. Finally, assumption W3 provides control over the relative tail behavior of the log-likelihood, requiring that log-likelihood ratios decay sufficiently fast—in an integrable manner—as the parameter value diverges, avoiding arbitrary peaks of the likelihood that may cause the MLE to overfit the data.
Walker, (1969) showed that assumptions W1–W3, together with a positive and continuous prior density, imply posterior consistency at any $\theta_{\star}\in\Theta$ as a consequence of MLE consistency. While technically valid—and while the positivity of the prior density may, under mild regularity conditions, be viewed as equivalent to the KL support condition of Schwartz, (1964)—these assumptions are quite specific and restrictive (see, e.g., Section 5 for a demonstration of this point). More importantly, their formulation is not informed by any clear connection to the core mechanisms that govern posterior consistency, but rather aim to obtain the latter as a consequence of well-behaved frequentist procedures. Before presenting our own alternative treatment, we briefly review a more recent approach to consistency which, although potentially more broadly applicable, shares many of the same underlying limitations when applied to parametric models.